Estimate the pH of the resulting solution prepared by mixing 1.0 mole of solid disodium phosphate (Na2HPO4) and 1.25 mole of hydrochloric acid. The pKas of phosphoric acid are 2.1, 7.2, 12.4. Please explain each step
Answer #1

The reaction occurs:

HPO4-2 + H + = H2PO4-

All moles of HPO4-2 are neutralized and begins to form:

H2PO4- + H + = H3PO4

Then you have:

n H3PO4 = 0.25 mol

n H2PO4- = 0.75 mol

You have the expression:

pH = pKa + log (n H2PO4- / n H3PO4) = 2.1 + log (0.75 / 0.25) = 2.56

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.